Follow the systematic procedure (9 steps, if applicable) to calculate for the
molar solubility of BaSO4 in a solution in which [H3O+] is 0.75 M.
1
Expert's answer
2020-10-12T13:32:17-0400

The equation of solubility is


"\text{BaSO}_4(s)\rightleftarrows\text{Ba}^{2+}(aq)+\text{SO}_4^{2-}(aq)."

The expression for solubility product is


"k=[\text{Ba}^{2+}][\text{SO}_4^{2-}]=1.08\cdot10^{-10}"

The equilibrium reactions and equilibrium constant equations:


"\text{H}_2\text{SO}_4(aq)+\text{H}_2\text{O}(l)\rightleftarrows\text{HSO}_4^-(aq)+\text{H}_3\text{O}^+(aq)\\\space\\

k_{a1}=\frac{[\text{HSO}_4^-][\text{H}_3\text{O}^+]}{[\text{H}_2\text{SO}_4]}=10^3,\\\space\\

\text{HSO}_4^-(aq)+\text{H}_2\text{O}(l)\rightleftarrows\text{SO}_4^{2-}(aq)+\text{H}_3\text{O}^+(aq)\\\space\\

k_{a2}=\frac{[\text{SO}_4^{2-}][\text{H}_3\text{O}^+]}{[\text{H}\text{SO}_4^-]}=0.0102."

Now we can substitute the last equation for ka2 in equation for ka1:

"[\text{H}_2\text{SO}_4]=\frac{[\text{SO}_4^{2-}][\text{H}_3\text{O}^+]}{0.0102\cdot10^3}"

Assume that the solubility is


"s=[\text{Ba}^{2+}]=[\text{SO}_4^{2-}]+[\text{H}\text{SO}_4^-]+[\text{H}_2\text{SO}_4]"

Substitute the previous equation and equations for ka2 and ka1, numerical values, and obtain expression for s:

"s=\sqrt{(1.08\cdot10^{-10})(1+98.039[\text{H}_3\text{O}]+0.98039[\text{H}_3\text{O}]^2)}."

For [H3O+] = 0.75 M the calculations give us


"s=8.975\cdot10^{-5}\text{ M}"